OpenStudy (faman39):

Which of the following would you use as the percent equation to determine that 70 is 40% of what number? 40 = 0.7 * a a = 0.7 * 40 a = 0.4 *70 70 = 0.4 *a

OpenStudy (y2o2):

a = 0.4 *70

OpenStudy (y2o2):

\[a = {{40\over 100} \times 70} = 0.4 \times 70\]

OpenStudy (faman39):

a = 0.7 * 40

OpenStudy (faman39):

Right?

OpenStudy (anonymous):

Look at the way I translate " 70 is 40% of what number? " 70 = .4 * a

OpenStudy (faman39):

70 = 0.4 *a

OpenStudy (faman39):

Thnk u all :) Thnks chloro! :)